AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 08:28:17 2025
Author
*
:
Subject
*
:
<p>07:00</p> <p>Run 12 started. D+ beam is incident on the target at 1.5 MeV</p> <p>Beam intensity in the ring is ~4e7 particles</p> <p>Target density shows some fluctuations but is ~1e11 atoms/cm2. Target temp 35K</p> <p>Detectors are positioned fully out (Left arm 18.86mm and right arm 19.8 mm ) and are not moving</p> <p>Ring pressure looks OK, YR09 E target sensor seems unaffected by the target currently reads ~5e-11 mbar.</p> <p>Beam is in the ring for ~15s.</p> <p>Temps and system checks OK</p> <p>Leakage current OK</p> <p>Merger and tape server OK</p> <p>Horizontal beam position 8 mm and vertical beam position 0mm for overlap with target.</p> <p>Statistics - attachment 1</p> <p>Energy histogram - attachment 2</p> <p>XY plot - attachment 3</p> <p>We see high rate at 1.5 MeV in left dssds but very little in right dssds. XY plots also show counts are concerntrated in a very circular distribution.</p> <p>07:15</p> <p>Gleb has just called and requires access to the cave to check on a turbo pump in the source.</p> <p>Ring vacuum looks good</p> <p>Beam will be down for 10-15 minutes only hopefully. </p> <p>07:45</p> <p>Beam is back in the ring but injection requires optimisation.</p> <p>09:00</p> <p>There are some issues with the source that require a few hours to resolve. Potentially until 7 or 8 pm tonight. We have about an hour before work starts.</p> <p>We have done a quick test of the motors and moved the left and right arms in using the pnuematic only.</p> <p>New position left arm = 62.71 mm , right arm = 64.6 mm.</p> <p>We observed the rates in left detectors went up significantly. In the right detectors we observed little increase.</p> <p>See attached xy plot - attachment 4</p> <p>See attached energy histogram - attachment 5</p> <p>Beam intensity has dropped significantly - below 1e6 particles so even at in positons we see few counts at the moment.</p> <p>09:25</p> <p>Source has degraded pretty rapidly so beam will be stopped for now.</p> <p>Detectors moved back out.</p> <p>Target nozzle moved from X= -13000, Y=2500 to X= -11000, Y=2500. S1 has dropped to 5e-10 mbar.</p> <p>Note Gleb mentioned that last night it was likely a large proportion of the ions stored were molecular hydrogen not deuterium.</p> <p>16:18</p> <p>Temps and system checks OK</p> <p>Leakage current OK</p> <p>Merger and tape server OK</p> <p> </p> <p>18:10</p> <p>Temps and system checks OK</p> <p>Leakage current OK</p> <p>Merger and tape server OK</p> <p> </p> <p>19:38</p> <p>Operators are trying to optimize some beam. For now the beam is lower than it should be probably because is easy to work with the source like that.</p> <p>Temps and system checks OK</p> <p>Leakage current OK</p> <p>Merger and tape server OK</p> <p><br /> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown